# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v4

@@ 2.3 Applications in Function Approximation and Multivariable Function Decomposition
3 Kolmogorov-Arnold Network Architecture
3.1 Basic Structure and Features of KAN Networks
The architecture of Kolmogorov-Arnold Networks (KAN) is unique in that its core idea is to replace traditional fixed linear weights with learnable univariate functions, achieving greater flexibility and adaptability. The architecture of KAN consists of multiple layers, each containing several nodes and edges. Each node is responsible for receiving input signals from the previous layer and applying nonlinear transformations to these signals via learnable univariate functions on the edges. These univariate functions are typically parameterized by spline functions to ensure smoothness and stability in the data processing process [5].
The advantage of this architecture is its ability to dynamically adapt to different data patterns, enhancing the model’s expressiveness. Specifically, each weight parameter in KAN is represented as a univariate function, adjusted during the training process to adapt to different data patterns and improve the network’s flexibility and accuracy [10, 5].
3.2 What is KAN
Univariate Function Representation
A core feature of KAN is that each weight parameter is represented by a learnable univariate function. These functions are typically implemented using spline functions or other suitable parameterization methods. This design allows KAN to apply dynamically adjusted activation functions on the network edges, rather than using fixed activation functions at the nodes as in traditional neural networks.
Dynamic Activation Functions
In KAN, activation functions are not limited to predefined forms (such as ReLU, sigmoid, etc.) but can be dynamically adjusted during the training process. This flexibility enables KAN to adapt to various complex data patterns and nonlinear relationships.
Architectural Flexibility
The architecture of KAN can be customized and adjusted according to the specific tasks and datasets. This flexibility is reflected in the representation of weight parameters and the dynamic adjustment of activation functions.
3.3 What KAN is Not
Fixed Weights and Activation Functions
Traditional neural networks, such as multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), use fixed linear weights and predefined activation functions. In these networks, the weight parameters are fixed during the training process and do not dynamically adjust.
Static Architectural Design
The architectural design of traditional neural networks is static and does not change during the training process. The forms of weights and activation functions are determined when the network is constructed.
3.4 Methods Unique to KAN
Weight Representation Based on Kolmogorov-Arnold Theorem
KAN uniquely employs the Kolmogorov-Arnold theorem, representing weight parameters through learnable univariate functions. This method allows KAN to capture more complex nonlinear relationships, which is not possible with traditional neural networks.
Dynamically Adjusted Activation Functions
KAN introduces dynamic activation functions, meaning the forms of activation functions can be adjusted based on data requirements during the training process. This flexibility significantly enhances the model’s expressiveness and adaptability.
Extensive Parameter Optimization
Unlike traditional neural networks that only optimize fixed weights and biases, KAN’s optimization process includes optimizing the parameters of activation functions. This multi-level optimization provides greater flexibility, allowing KAN to more accurately capture the complex relationships in data.
3.5 Specific Technical Optimizations
Residual Activation Functions
Using base functions $b(x)$ (similar to residual connections), where $b(x)=\text{silu}(x)=\frac{x}{1+e^{-x}}$, making the activation function $\Phi(x)$ a sum of the base function and spline functions.
Initialization Method
Each activation function is initialized as $spline(x)\approx 0$ (by setting $c_{i}\sim N(0,\sigma^{2})$, typically $\sigma=0.1$), while $w$ is initialized in the same way as the linear layers in MLPs.
Updating the Spline Grid
Dynamically update each grid based on input activation to address the problem of splines being defined on a bounded region while activation values may exceed this fixed region during training.
3.6 Is Deep Learning a Necessary Component of KAN?
The design philosophy of Kolmogorov-Arnold Networks (KAN) is rooted in the Kolmogorov-Arnold theorem, emphasizing the representation of weight parameters through univariate functions. This architectural design is compatible with the fundamental ideas of deep learning, but whether KAN requires all components of deep learning and the relationship between the two needs further exploration.
Necessity
Training Process
KAN relies on deep learning training methods such as gradient descent and backpropagation. These methods are indispensable in KAN’s optimization process, making the training process of deep learning essential for KAN. These optimization algorithms have proven effective in adjusting a large number of parameters, achieving good generalization performance.
Nonlinear Representation Capability
The multi-layer network architecture and nonlinear activation functions in deep learning provide strong nonlinear representation capabilities, consistent with KAN’s goal of capturing complex relationships through dynamically adjusted activation functions. Therefore, the core ideas of deep learning are irreplaceable in enhancing KAN’s performance. Through the multi-layer structure, deep learning can extract abstract features of data layer by layer, which is crucial in handling complex data.
Scalability and Adaptability
The scalability and adaptability of deep learning models are important aspects that KAN can leverage. Deep learning models have demonstrated strong adaptability and scalability in various fields such as computer vision, natural language processing, and reinforcement learning, and these experiences can be directly applied to the research and development of KAN.
Alternatives
Traditional Machine Learning Methods
Although KAN can leverage deep learning training methods, its unique representation of weight parameters through univariate functions and the design of dynamic activation functions differentiate it from traditional deep learning models. Theoretically, other optimization methods (such as evolutionary algorithms or Bayesian optimization) could be used to train KAN, but these methods might not fully realize KAN’s potential. Traditional methods might be less efficient in handling high-dimensional and complex datasets.
Other Architectures
If deep learning methods are not used to train KAN, new algorithms might need to be developed to address its complex optimization problems. However, the effectiveness and efficiency of such methods in practical applications remain to be verified. For example, symbolic regression or genetic algorithms could be used for similar tasks, but their scalability and optimization efficiency might not match that of deep learning techniques.
3.7 Is KAN a Branch of Deep Learning?
Theoretical Foundation
The theoretical foundation of KAN is the Kolmogorov-Arnold theorem, which differs from the multi-layer nonlinear network models of deep learning. However, both aim to capture complex nonlinear relationships through hierarchical structures. Deep learning models achieve this through architectures like multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), while KAN does so through dynamically adjusted activation functions and univariate function weights.
Architectural Features
Although KAN features unique univariate function weight representations, its hierarchical network structure and dynamic activation function design share similarities with neural network models in deep learning. Therefore, KAN can be seen as a special branch of deep learning, aiming to enhance model flexibility and expressiveness through new methods.
Application Domains
KAN’s application domains include data fitting, partial differential equation solving, hyperspectral image classification, and time series analysis, which are traditionally strengths of deep learning. KAN demonstrates the potential to surpass traditional deep learning models in these areas through its unique architecture and optimization methods.
Figure 1: Example of KAN’s 655Application
3.8 Position and Role of Activation Functions in KAN Networks
One of the key features of KAN is the use of learnable activation functions. These activation functions are typically constructed from spline functions, a type of piecewise polynomial function defined by a set of control points that determine their shape. These control points can be adjusted independently, allowing local changes in the function’s shape without affecting its global properties. This characteristic ensures that spline functions maintain the smoothness of the activation functions while flexibly adapting to complex patterns in the data [10].
By using spline functions, KAN is able to apply complex nonlinear transformations on each edge, enhancing the network’s performance in handling nonlinear relationships. Specifically, the diversity and flexibility of spline functions ensure that the model remains efficient and stable when dealing with different types of data [5].
3.9 Function Approximation Capability and Neural Scaling Effect of KAN Networks
Kolmogorov-Arnold Networks (KAN) enhance their ability to handle nonlinear relationships by using spline functions to implement complex nonlinear transformations on each edge. The diversity and flexibility of spline functions ensure that the model remains efficient and stable when dealing with various types of data. KAN’s training process adopts a backpropagation algorithm similar to that of traditional neural networks, but the core difference lies in simultaneously optimizing the learnable activation functions on the edges and the weights between nodes. This requires a fine-tuned adjustment strategy to ensure model stability and convergence.
To improve training efficiency and generalization ability, research suggests using regularization techniques such as Dropout, weight decay, and batch normalization. Additionally, combining advanced optimization algorithms like Adam and stochastic gradient descent can further optimize KAN’s performance when handling large-scale datasets [5].
The scaling perspective of KAN can be summarized as follows:
  1. 1.
    Simple two-layer KANs may provide non-smooth representations, while deeper KANs can achieve smoother function approximation by increasing the number of layers.
  2. 2.
    Even KANs with a limited number of dimensions can effectively approximate target functions, demonstrating the broad approximation capability of KAN.
The neural scaling law describes the phenomenon where model test loss decreases with an increase in the number of model parameters, following the relationship $\ell\propto N^{-\alpha}$, where $\ell$ represents the root mean square error (RMSE) during testing, $N$ represents the number of parameters, and $\alpha$ is the scaling exponent. This indicates that increasing the number of parameters can significantly improve model accuracy.
KAN leverages the precise approximation capability of B-spline curves by fine-graining them over the data domain, achieving more accurate function representation. This strategy shows the unique advantage of KAN compared to multilayer perceptrons (MLPs), which typically do not train on data partitioning. By starting with a KAN with fewer parameters and gradually refining the spline grid to a model with more parameters, KAN can achieve high-precision approximation with lower complexity.
@@ 5.3 Methods to Ensure Stability and Efficiency of the Training Process
6 Applications of KAN Networks
6.1 Symbolic Regression
Symbolic Regression is a method that seeks mathematical formulas or expressions that best represent the data. Unlike traditional regression analysis, symbolic regression not only fits data points but also attempts to discover the underlying mathematical relationships that generate the data. This method is particularly important in scientific research because it can reveal the physical laws and mathematical principles behind the data.
Despite its theoretical potential, symbolic regression faces many challenges in practical applications. Firstly, symbolic regression requires searching for the best mathematical expression in a vast space of formulas, which is computationally intensive. Traditional symbolic regression algorithms, such as those based on genetic programming, can generate highly interpretable models but are often limited by computational resources, making it difficult to handle large-scale datasets and high-dimensional data [15].
Secondly, symbolic regression performs poorly when dealing with noisy data. Real-world data often contains various noise and outliers, making it more difficult to accurately extract the underlying mathematical relationships. Traditional methods tend to overfit or fail to find robust formulas under such circumstances [16].
6.1.1 Advantages of KAN in Symbolic Regression
Kolmogorov-Arnold Networks (KAN) exhibit significant advantages in the application of symbolic regression, particularly in addressing the aforementioned challenges. KAN’s unique architecture and flexible activation functions enable it to effectively perform symbolic regression tasks, as demonstrated below:
High Precision and Interpretability
KAN uses spline functions as activation functions, allowing for more precise fitting of complex nonlinear relationships. Compared to traditional neural networks, KAN can dynamically adjust its activation functions to better capture complex patterns in the data. For example, Liu et al. (2024) demonstrated KAN’s superior performance in symbolic regression, effectively modeling complex data and achieving high-precision formula discovery [5].
Innovations and Improvements
Path Signature Integration: Path signatures are tools that capture the geometric features of time series data. By integrating path signatures, KAN can better understand and predict complex time series patterns. Path signatures effectively capture the geometric shape and structural information of data, providing a more comprehensive perspective for modeling time series data [17].
Dynamic Adaptation: KAN can dynamically adapt to different data patterns by learning adjustable parameters of path signatures, improving its performance in multi-task time series forecasting. The introduction of path signatures allows KAN to dynamically adjust according to different time series characteristics, enhancing the model’s adaptability and robustness [18].
KAN 2.0: Enhancing Symbolic Regression
Recent advancements in the field have further bolstered KAN’s capabilities in symbolic regression. Liu et al. (2024), in their study titled "KAN 2.0: Kolmogorov-Arnold Networks Meet Science," introduced several new tools such as MultKAN and the KAN compiler. These tools allow KAN to perform symbolic regression with even greater precision, especially in scientific contexts where uncovering underlying mathematical relationships is crucial. The dynamic adjustment capabilities provided by these tools enable KAN to better model and discover complex symbolic formulas, making it an invaluable asset in the realm of scientific discovery[19].
The emergence of KAN has had a profound impact on the field of symbolic regression. Firstly, the high precision and interpretability of KAN allow researchers to better understand the mathematical relationships behind the data, thereby advancing scientific discovery. For example, Kundu et al. (2024) demonstrated KAN’s superior performance in symbolic regression, showing through experiments that KAN is more accurate and efficient than traditional methods when handling complex datasets. Secondly, the flexibility and robustness of KAN make it particularly effective in dealing with noisy data. By dynamically adjusting activation functions, KAN can accurately identify the symbolic formulas that generate the data even in high-noise environments, providing higher reliability and stability for practical applications [5].
6.2 Time Series Prediction

### v5

@@ 4.3 Methods to Ensure Stability and Efficiency of the Training Process
5 Comparison between MLP and KAN
The Multi-Layer Perceptron (MLP) is one of the most fundamental structures in the field of deep learning. In this section, we will compare MLP and KAN in several aspect.Since KAN requires incorporating an external knowledge base, it results in a more complex time complexity[15]. In general, KANs exhibit potential across various fields; however, they consume more computational resources.
5.1 Performance on Noisy Functions
KAN seems to be a promising way to deal with data with noise. Zen et al [16] compare KAN and MLP based on 6 different irregular or nosy functions. Their research shows that, for some function, KAN has a better performance. However, on regular function noise, KAN performs worse than MLP, even though KAN converges more faster. For jumping function, MLP performs better.Shukla et al [17] compare KAN-based architectures and MLP for solving PDEs with noisy data. The result shows a similar tendency. They find that KAN has a better performance than MLP after specific parameter tuning.
5.2 Time Complexity
KAN may take more time compared with MLP in the same task.After [16] comparing the time consumption by L-BFGS algorithm, Zen et al find that KAN spends more time in almost all task. Moreover, KAN even takes 100 times the time of MLP, which extremely limit its application to the real world. In the experiment done by Sasse and Farias [18], they use KAN based on B-spline activation function (Spline-KAN) for Federated Learning (FL) tasks. Their result show that although KAN does spends more time than MLP, which may take 1.36 times the time of MLP. However, this difference is not as significant as expected.
5.3 Low-Data Training
KAN shows a worse performance than MLP when focusing on low-data regimes. Pourkamali-Anaraki [19] assumes that MLP has advantages than KAN when using limited labeled data after comparing the learnable parameters of the two. In general, KAN has more learned parameters. Zen et al [16] finds that after adding data, KAN will perform better in some specific tasks. However, it shows that MLP can be more accurate than KAN if the training is limited.
5.4 Reinforcement Learning
Conservative Q-learning (CQL) task comparison done by Guo et al[15] shows that pure MLP-based CQL model has the best performance. However, adding more hidden layers to the KAN actor did not significantly improve the model’s performance.
5.5 Performance on Computer Vision
Cheon [20] validated the effectiveness of KAN in visual tasks on the MNIST, CIFAR10, and CIFAR100 datasets. The result shows that KAN has a better performance than MLP on CIFAR10, and CIFAR100 but slightly worse than the state-of-the-art ResNet-18. Bodner et al [20] create Convolutional Kolmogorov-Arnold Networks (Convolutional KANs) and validate the performance of Convolutional KANs against traditional architectures across Fashion-MNIST dataset. Their experiments show that Convolutional KANs reach the a similar level of accuracy with half parameters which demonstrate the potential of KANs in computer version field. However, Tran et al [21] explore the limits of KAN on classification tasks. They claim that KAN exceed MLP despite theoretical advantages in accuracy and interpretability since KAN will consume more resources and have higher latency than MLPs. Their results show that MLP is still the better option when it comes to classification task.
5.6 Performance on Tabular Data
Poeta et al [22] postpone a benchmarking study comparing MLP and KAN on tabular datasets. The results show that KANs have a higher F1 score and more accuracy, demonstrating robust in processing complex data. However, they also highlight the high cost of KAN.
5.7 Performance on Time Series Problems
Vaca-Rubio et al [23] compare KAN and MLP in a real-world satellite traffic forecasting task. Their results show that KAN has less learnable parameters. The reduction in complexity shows that KAN can achieve higher accuracy with simpler and lighter structure.
Performance on Dynamical Systems
Nehma and Tiwari [24] compare MLP and KANs in finding Deep Koopman operators for linearizing nonlinear dynamics. In their work, they construct a KANs-based deep Koopman framework and apply to the pendulum, the combined pendulum-cart system and an orbital Two-Body Problem (2BP) for data-driven discovery of linear system dynamics. Their results show that KANs is superior to MLP DNNs, learning 31 times faster, being 15 times more parameter-efficient, and predicting 1.25 times more accurately in the 2BP case. They claim KANs can help develop Deep Koopman Theory.
6 Application of KAN Networks
In this section, we will discuss the applications of KAN in different fields.
6.1 Time Series